# When Large Language Models Meet Vector Databases: A Survey

### v3

@@ III-B VecDB as a Cost-effective Semantic Cache
One of the primary benefits of integrating VecDBs with LLMs is the significant reduction in data operational costs [35]. GPT-Cache [36], a VecDB that serves as a semantic cache, as shown in Figure 2, for instance, stores responses to previously asked queries and works as a cache before calling LLM APIs. This caching mechanism means that the system doesn’t need to have API calls to wait for generated responses from scratch every time, reducing the number of costly API calls to the LLM. Moreover, this approach also speeds up the response time, enhancing the user experience.
@@
III-C VecDB as a Reliable Memory of GPTs
@@
To solve this problem, applications generally use VecDBs that serve as a robust memory layer [37] that addresses one of the intrinsic limitations of LLMs: the static nature of their knowledge. While LLMs excel in generating human-like text based on patterns learned during training, they cannot inherently update their knowledge base dynamically and, thus, may lack the few-shot learning ability. VecDBs bridge this gap by offering a storage solution that can be continually updated with new information, ensuring that the LLM’s responses are informed by the most current and relevant data available.
@@
IV Expanding Horizons: RAG Advancements
IV-A Multimodality of RAG
@@
RAG has now evolved to handle a wide range of data types by lending the power of multimodal models. The impressive achievements of LLMs have inspired significant advancements in vision-language research. DALL-E from OpenAI introduced a Transformer-based approach for converting text to images, treating images as sequences of discrete tokens. Subsequent improvements in the text-to-image area  [38] have been achieved through methods like model scaling, pre-training, and enhanced image quantization models. BLIP-2  [39] uses static image encoders with LLMs for efficient visual language pre-training, facilitating direct image-to-text transformations. Flamingo  [40] presented a visual language model for text generation, showcasing remarkable adaptability and leading performance across various vision-and-language tasks. CM3  [41] trained a randomly masked model on a large HTML corpus and showed that the model is capable of generating images and text. FROMAGe  [42] gains robust multimodal capabilities for few-shot learning solely from image-caption pairs, unlike other models that necessitate large-scale, interwoven image-text data from the websites.
@@
To import speech data to RAG systems, Wav2Seq  [43] allows for efficient pre-training without the need for transcriptions, using techniques like k-means clustering and byte-pair encoding to generate pseudo-subwords from speech. The Large Language and Speech Model (LLaSM)  [44] is an end-to-end trained, large multi-modal speech-language model equipped with cross-modal conversational skills and proficient in understanding and responding to combined speech-and-language directives. Videos are also made available for certain types of RAG systems. Vid2Seq  [45] enhances language models with specific temporal indicators for predicting event limits and textual descriptions in a single output sequence.
IV-B Retrieval Optimizations of RAG
@@
To better harness the knowledge from various types of data, kNN-LMs  [46] explore how incorporating nearest neighbor search into language models can enhance their ability to generalize by effectively leveraging memorized examples. EASE  [47] is distinctive in its use of entities as a strong indicator of text semantics, providing rich training signals for sentence embedding. In-context RALM  [48] proves that with the language model architecture unchanged, simply appending grounding documents to the input will improve the performance. SURGE  [49] enhances dialogue generation by incorporating context-relevant sub-graphs from a knowledge graph. Another work that combines knowledge graphs and LLMs is RET-LLM  [50], which is designed to equip LLMs with a general write-read memory unit. These studies have focused on retrieval granularity and data structuring levels, with coarse granularity providing more, but less precise, information. Conversely, structured text retrieval offers detailed information at the cost of efficiency.
@@
To utilize both internal knowledge and external resources, SKR  [51] improves LLMs’ ability by enabling them to assess what they know and identify when to seek external information to answer questions more effectively. Selfmem  [52] enhances retrieval-augmented text generation by creating an unbounded memory pool using the model’s output and selecting the best output as memory for subsequent rounds. FLARE  [53] uses predictions of upcoming sentences to anticipate future content and retrieve relevant documents for regenerating sentences, especially when they contain low-confidence tokens. Atlas  [54] demonstrates impressive performance in tasks like question-answering and fact-checking, outperforming much larger models despite having fewer parameters. Many other works like these also aim to make the RAG system more efficient and competent.
V Discussion: Challenges and Future Work
V-1 Are vector searches all you need?
@@
Although vector data is an efficient representation of unstructured data, vector-based search methodologies are still not well suited for operations such as structured queries (SQL), post-query filtering, comprehensive full-text searches, relation-based graph, and keyword search mechanisms that are fundamental to conventional database systems and search engines. In the context of diverse applications for different users with multiple needs, vector search, or any single search approach, is not a one-size-fits-all solution. For example, for searching exact transaction data, users may have queries of very specific keywords to look up in a huge pool of data; giving them semantic approximate results totally would not make sense, resulting in a one-step forward yet two-step back. On the other hand, traditional DBMSs are well-developed for efficiently handling structured data within their transactional and analytical applications, and knowledge graph can represent the interation and logic between events and entities [55], in both cases, a dedicated VecDB is not yet ready for those queries.
@@
V-2 VDBMS with multi-modality
In real-world data representation and search, we may look for multi-modal data entries that are not restricted to a specific kind of structured, semi-structured, or unstructured data, but with any random combination of them, e.g., users can look for a knowledge graph from an interacted prompt of image and description. While searches using uni-modal data may lack the prospect of providing informative and more contextually appropriate search results, multi-modal data search and its hybrid processing also present a non-trivial challenge for VecDBs’ storage and retrieval [56], since integrating and merging multi-modal data require efficient preprocessing with multi-modal encoders, multi-modal storage indexes, but also ranking weight assignment and multi-modal data fusion.
VI Conclusion

### v4

@@ 3.2 VecDB as a Cost-effective Semantic Cache
One of the primary benefits of integrating VecDBs with LLMs is the significant reduction in data operational costs Sanca and Ailamaki (2023). GPT-Cache Bang (2023), a VecDB that serves as a semantic cache, as shown in Figure 2, for instance, stores responses to previously asked queries and works as a cache before calling LLM APIs. This caching mechanism means that the system doesn’t need to have API calls to wait for generated responses from scratch every time, reducing the number of costly API calls to the LLM. Moreover, this approach also speeds up the response time, enhancing the user experience.
@@
3.3 VecDB as a Reliable Memory of GPTs
@@
VecDBs serve as a robust memory layer Zhang et al. (2023b) that addresses one of the intrinsic limitations of LLMs: the static nature of their knowledge. While LLMs excel in generating human-like text based on patterns learned during training, they cannot inherently update their knowledge base dynamically and, thus, may lack the few-shot learning ability. VecDBs bridge this gap by offering a storage solution that can be continually updated with new information, ensuring that the LLM’s responses are informed by the most current and relevant data available.
@@
4 Expanding Horizons: RAG Advancements
4.1 Multimodality of RAG
@@
RAG has now evolved to handle a wide range of data types by lending the power of multimodal models. The impressive achievements of LLMs have inspired significant advancements in vision-language research. DALL-E from OpenAI introduced a Transformer-based approach for converting text to images, treating images as sequences of discrete tokens. Subsequent improvements in the text-to-image area  Zhang et al. (2023a) have been achieved through methods like model scaling, pre-training, and enhanced image quantization models. BLIP-2  Li et al. (2023b) uses static image encoders with LLMs for efficient visual language pre-training, facilitating direct image-to-text transformations. Flamingo  Alayrac et al. (2022) presented a visual language model for text generation, showcasing remarkable adaptability and leading performance across various vision-and-language tasks. CM3  Aghajanyan et al. (2022) trained a randomly masked model on a large HTML corpus and showed that the model is capable of generating images and text. FROMAGe  Koh et al. (2023) gains robust multimodal capabilities for few-shot learning solely from image-caption pairs, unlike other models that necessitate large-scale, interwoven image-text data from the websites.
@@
To import speech data to RAG systems, Wav2Seq  Wu et al. (2022) allows for efficient pre-training without the need for transcriptions, using techniques like k-means clustering and byte-pair encoding to generate pseudo-subwords from speech. The Large Language and Speech Model (LLaSM)  Shu et al. (2023) is an end-to-end trained, large multi-modal speech-language model equipped with cross-modal conversational skills and proficient in understanding and responding to combined speech-and-language directives. Videos are also made available for certain types of RAG systems. Vid2Seq  Yang et al. (2023) enhances language models with specific temporal indicators for predicting event limits and textual descriptions in a single output sequence.
4.2 Retrieval Optimizations of RAG
@@
To better harness the knowledge from various types of data, kNN-LMs  Khandelwal et al. (2020) explore how incorporating nearest neighbor search into language models can enhance their ability to generalize by effectively leveraging memorized examples. EASE  Nishikawa et al. (2022) is distinctive in its use of entities as a strong indicator of text semantics, providing rich training signals for sentence embedding. In-context RALM  Ram et al. (2023) proves that with the language model architecture unchanged, simply appending grounding documents to the input will improve the performance. SURGE  Kang and Kwak (2023) enhances dialogue generation by incorporating context-relevant sub-graphs from a knowledge graph. Another work that combines knowledge graphs and LLMs is RET-LLM  Modarressi et al. (2023), which is designed to equip LLMs with a general write-read memory unit. These studies have focused on retrieval granularity and data structuring levels, with coarse granularity providing more, but less precise, information. Conversely, structured text retrieval offers detailed information at the cost of efficiency.
@@
To utilize both internal knowledge and external resources, SKR  Yu et al. (2023) improves LLMs’ ability by enabling them to assess what they know and identify when to seek external information to answer questions more effectively. Selfmem  Xin et al. (2023) enhances retrieval-augmented text generation by creating an unbounded memory pool using the model’s output and selecting the best output as memory for subsequent rounds. FLARE  Jiang et al. (2023) uses predictions of upcoming sentences to anticipate future content and retrieve relevant documents for regenerating sentences, especially when they contain low-confidence tokens. Atlas  Izacard et al. (2022) demonstrates impressive performance in tasks like question-answering and fact-checking, outperforming much larger models despite having fewer parameters. Many other works like these also aim to make the RAG system more efficient and competent.
5 Discussion: Challenges and Future Work
5.0.1 Are vector searches all you need?
@@
Although VecDBs offer a cost-efficient modality for information retrieval within LLMs’ frameworks, their utility in traditional relational database operations remains limited. Specifically, vector-based search methodologies are still not well optimized for operations such as post-query filtering, comprehensive full-text searches, and keyword search mechanisms that are fundamental to conventional database systems and search engines. In the context of diverse applications for different users with multiple needs, vector search, or any single search approach, is not a one-size-fits-all solution. For example, for searching exact transaction data, users may have queries of very specific keywords to look up in a huge pool of data; giving them semantic approximate results totally would not make sense, resulting in a one-step forward yet two-step back. On the other hand, traditional DBMSs are well-developed for efficiently handling structured data within their transactional and analytical applications, while dedicated VecDBs are not yet.
@@
5.0.2 VDBMS with multi-modality
In real-world data representation and search, we may look for multi-modal data entries that are not restricted to a specific kind of structured, semi-structured, or unstructured data, but with any random combination of them, e.g., users can look for a knowledge graph from an interacted prompt of image and description. While searches using uni-modal data may lack the prospect of providing informative and more contextually appropriate search results, multi-modal data search and its hybrid processing also present a non-trivial challenge for VecDBs’ storage and retrieval Wang and et al. (2023), since integrating and merging multi-modal data require efficient preprocessing with multi-modal encoders, multi-modal storage indexes, but also ranking weight assignment and multi-modal data fusion.
5.0.3 Data preprocessing
While text embedding is considered efficient for processing long text and its retrieval, it comes with a dazzling challenge for building every text-based knowledge database. It is observed that vector retrieval would have fundamentally better performance when precise chunks of text with clear meaning are applied. This raises an interesting question: how do we have a proper (probably unified) embedding methodology for every raw text?
On the other hand, dimension reduction for vector data is considered important since compressing high-dimensional vectors will use less memory and disk resources. From this view, the current vector data dimension is not cost-effective, as the vectors are large, e.g., a 1536-dimensional vector embedded only by OpenAI’s lightweight model text-embedding-3-small  [^6] https://platform.openai.com/docs/guides/embeddings is about 6 kilobytes; with scaling up to 1 billion, the data size would be 6 terabytes. Also with the curse of dimensionality theory, data sparsity increases with dimensionality, resulting in hardship in distinguishing meaningful similarities. Moreover, current dimension reduction algorithms like PQ contribute to a memory-efficient vector data representation, whereas lossy data compression still hurts the recall quality Chen et al. (2021). This requires a more efficient vector data dimension reduction algorithm for VecDBs and searches.
5.0.4 Data management systems for LLMs with multi-tenancy
Both for VecDBs and LLMs, this challenge is particularly in preventing data leaks between tenants and maintaining the isolation and security of distinct tenants’ data while ensuring efficiency, where isolation is crucial for privacy and security. On the other hand, it must be balanced with the need for dynamic resource sharing to ensure resources are flexibly allocated and deallocated.
5.0.5 Cost-effective and scalable data storage and retrieval
While state-of-the-art vector search algorithms are mainly memory-based, memory resources in the cloud are considered expensive and volatile. With billion-scale amounts of data, naively handling them in memory resources is bankrupting. On the other hand, disk-based search methods have downgraded performance that does not satisfy real-time applications with low latency tolerance. This calls for a more cost-effective way to ensure the performance of LLM applications.
5.0.6 Knowledge conflict
Conflicting knowledge from the same or different knowledge bases presents a non-trivial challenge for both humans and LLMs to distinguish the correct piece of knowledge. This conflict becomes particularly prominent when integrating multiple knowledge bases where each potentially carries its own biases or inaccuracies, such as when different datasets provide contradictory information about the same fact or when biases inherent in data sources lead to inconsistent conclusions. While VecDBs cannot naturally distinguish, resolving such conflicts requires robust conflict resolution strategies that can assess the reliability of sources and the context of data to determine the most accurate information.
6 Conclusion